OpenStudy (anonymous):

Please check my work! Solving algebraic identities. (y+x)(ax+b) yax + yb + ax^2 + xb y^2 + a^2 + x^4 + b^2 @hero @dan815 @Abhisar @pooja195@nincompoop @Mehek14 @jigglypuff314

OpenStudy (anonymous):

Please I really need help @Jhannybean @jigglypuff314

OpenStudy (anonymous):

Please check my work @dan815 @imqwerty

OpenStudy (anonymous):

@DLS @whpalmer4 @Mehek14

OpenStudy (anonymous):

@Hero @nincompoop @nightcore144

OpenStudy (anonymous):

PLEASE HELP! Really... @GracieBugg @whpalmer4 @tHe_FiZiCx99

OpenStudy (anonymous):

@ryanvarghese12790 @Elsa213 @YoungStudier

OpenStudy (anonymous):

@undeadknight26 @imqwerty @Ondinana

OpenStudy (anonymous):

@praxer @aaronq @sleepyjess

OpenStudy (anonymous):

(y+x)(ax+b) yax + yb + ax^2 + xb y^2 + a^2 + x^4 + b^2 Is this right?

OpenStudy (anonymous):

@Ghostgate

OpenStudy (whpalmer4):

\[(y+x)(ax+b) = y(ax+b) + x(ax+b)\]by distributive property \[y(ax+b) + x(ax+b) = axy + by + ax^2 + bx\]again by distributive property

OpenStudy (whpalmer4):

I am not sure how you got from your correct 2nd line to your incorrect 3rd line...